OpenStudy (anonymous):

Find the 11th term of the arithmetic sequence. a3= 20.5 and a8= 13,a11=

OpenStudy (anonymous):

an = a1 + (n-1)d use a3 and a8 to find a1 and d then find a11

OpenStudy (anonymous):

so which is n

OpenStudy (anonymous):

a3 here n=3

OpenStudy (anonymous):

a8 here n=8 and a11 here n=11

OpenStudy (anonymous):

so then those are the answers

OpenStudy (anonymous):

no

OpenStudy (anonymous):

cuz when i put the 3 in i just got 3

OpenStudy (anonymous):

a3=a1+(3-1)d 20.5 = a1 + 2d ....i a8 = a1 + (8-1)d 13 = a1 + 7d ....ii

OpenStudy (anonymous):

can u solve equation i and ii to find a1 and d?

OpenStudy (anonymous):

no they didnt explain that

OpenStudy (anonymous):

subtract equation ii from i and we get, 20.5 -13 =a1 +2d -a1 -7d 7.5 = -5d d= -7.5/5 d=-1.5 Now, when d= -1.5 then from equation i , 20.5 = a1 + 2*(-1.5) a1=23.5 Now, a11 = a1 + 10d = 23.5 + 10*(-1.5) =8.5

OpenStudy (anonymous):

got it???

OpenStudy (anonymous):

sry no

OpenStudy (anonymous):

do u know that nth term of a arithemetic sequence is an = a1 + (n-1)d

OpenStudy (anonymous):

i know that but they never said anything about i

OpenStudy (anonymous):

well, you should know it before trying to solve these kind of problems

OpenStudy (anonymous):

thier no other way to solve but this way

OpenStudy (anonymous):

yes there are but all uses the idea an=a1+(n-1)d

OpenStudy (anonymous):

i noe how to use the formula but what confushes me is the I

OpenStudy (anonymous):

???

OpenStudy (anonymous):

nvr so i went back to what ur talking about is 9.27?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

what the the problem